# Affordable Generative Agents

### v1

@@ 4.3 Evaluation Protocol
We focus on token consumption and performance. We carry out a series of experiments to calculate the mean and standard deviation of token consumption. To comprehensively evaluate the agent’s behavior, we first reuse the evaluation methods of Generative Agents excluding manual assessments. These methods encompass interviewing agents with manually designed questions and conducting case study-specific activities.
Moreover, we propose two novel methods to evaluate the emergent social behaviors of generative agents:
@@
  • LLM Evaluator. Chiang & Lee (2023) have validated LLM as an evaluator, demonstrating comparable performance to expert evaluators. Consequently, we employ GPT-4 to evaluate the activities and dialogues generated by LLM agents. A 5-point Likert scale questionnaire is deployed to differentiate whether these activities and dialogues are from human or AI-generated.
@@ 5.1 Results on Generative Agents
Figure 3 illustrates the ablation study of token consumption based on multiple two game day simulations of 3-person town. Compared to the baseline, the cost of using only the Lifestyle Policy is 40.2% of the original, while using only the Social Relationship Memory is 58.6%, and the full AGA framework is 31.1%. It should be noted that the advantage of the Lifestyle Policy lies in the reuse of existing LLM-generated strategies. Hence, both Lifestyle Policy and AGA conduct experimental comparisons after running multiple iterations, and then storing a certain number of policies.
We also present the results of different configuration simulations in Table 1, where the cost for a 25-person town is 42.7% of the original. In the 25-person town, the probability of interactions between agents increases with the rise in population density. After triggering a dialogue, an agent adjusts its plan, invoking the reactions module to handle unexpected events. The occurrence of more events leads to more frequent output of reflections for deeper thinking. In order to ensure the flexibility of the LLM, we did not modify the reactions and reflections modules of the Generative Agents, which incurs additional costs for the 25-person version.
Case Study
We employ the same method used for Generative Agents to evaluate believable agents in this part. We conduct interviews with agents to evaluate their ability to remember past events, strategize for future tasks based on those memories, react appropriately to unexpected situations, and reflect on their previous actions to improve their future performance. We use the same questions as Generative Agents to inquire about five aspects: self-knowledge, memory, plans, reactions and reflections. Our findings indicate that our agents perform on par with the baseline, providing specific and accurate answers to the questions. Sample responses are provided in Appendix A.
In addition, we can validate the emergence of social behavior through specific events. In the 25-person town, Isabella Rodriguez, the owner of the coffee shop, plans to host a Valentine’s Day party and invite as many people as possible the following day, while Sam is running for mayor. After several experiments, we observe AGA got a similar result as baseline that about $12\sim 17$ persons attend the party and $4\sim 12$ persons knew Sam’s mayoral candidacy.
Relationship Evolution
In Social Memory, all initial relationships between agents are set to “Unknown” and subsequently updated following each interaction. We assess the emergence of social behavior by monitoring the relationship evolution. For example, in the 3-person town, Klaus Mueller and Maria Lopez discuss their research at a cafe, establishing a colleague relationship. Isabella Rodriguez, the cafe owner, engages in conversation while serving Klaus and Maria, transitioning to an acquaintance. A specific demonstration of social behavior is that agents modify their relationships with other agents through social interactions, thereby strengthening their ties with the community. We also conduct experiments on the 25-person environment. For visualization purposes, we quantify relationships on a scale of $0\sim 10$ (the higher value, the closer relationship. e.g., 0-strangers, 3-acquaintance, 5-co-worker, and 10-married). The result is shown in Figure 4. The data map reflects the actual relationships between agents. For example, in the virtual town, John Lin (JL in Figure 4) is profiled as living with his wife, Mei Lin (ML), and son, Eddy Lin (EL). Therefore, their relationship scores are all a perfect 10. Additionally, the data map shows changing relationships between agents, such as the relationship score between Tom Moreno (TM) and Yuriko Yamamoto (YY), who do not know each other, shifting to a 5. After conducting multiple experiments and calculating average scores, the data suggests an evolution towards an acquaintance-based community.
LLM Evaluator
The activities and dialogues conducted by the agents are recorded and evaluated using GPT-4, employing a 5-point Likert scale to discern whether the responses originated from an AI or a human. In this scale, a higher score indicates a closer resemblance to human-like responses. The complete questionnaire is provided in Appendix B.1. The results presented in Table 2 suggest that our method achieves scores comparable to the baseline. Due to the advantages of LLM in natural language dialogues, both methods approach human-level performance. However, the performance of agents in activity is considered potentially indicative of either AI or human behavior. The primary reason for suspecting AI is the overly detailed, minute-level activities generated by Generative Agents, which seems more akin to AI generation. A more comprehensive summary of the criteria used by GPT-4 to discern between AI and human entities can be found in Appendix B.2. We hope these suggestions from the LLM evaluator will guide us to create more believable agents in the future work.
@@
Table 1: The token cost with varying population setting
@@ 5.2 Results on VirtualHome
In the VirtualHome, we have designed an agent with the aim of experiencing a fulfilling day at home. The agent is required to generate a sequence of activities to achieve the goal. For each activity, the agent examines the items involved and the associated actions, creates a corresponding action sequence, and ultimately translates this sequence into actionable instructions, such as “<char0> [walk] <curtains> (32)”. The complete implementation is provided in the appendix C and an example activity generated by AGA is provided in Figure 5.
In contrast to Singh et al. (2023) and Huang et al. (2022) tested on manually designed household tasks, we aim for the agent to generate believable activities and decompose them into appropriate action sequences, leading to the absence of a clear method to determine task completion. Therefore, we request LLM agent to assess task completion based on the action sequences performed and the relevant objects, and decide whether to proceed to the next activity.
Table 3 presents the token consumption and task success rates for two methods on VirtualHome. The token data is derived from the total tokens required by the LLM agent to complete all activities in a day. $S_{LLM}$ refers to task success rate evaluated by LLM, while $S_{Human}$ means evaluated by human. Both evaluation methods indicate that AGA does not result in significant performance changes. However, AGA costs only 3.4% of the baseline. Because VirtualHome involves only single-agent interaction, the contribution predominantly stems from the Lifestyle Policy. In some precise tasks, LLM agents tend to predict task completion earlier, resulting in a higher success rate than human evaluations. Furthermore, our analysis revealed that the main cause of task failure is the inability to execute the generated plan with the items present in the current room or the available actions.
@@
Table 3: Cost and performance analysis in VirtualHome.
@@ 5.3 Further Analysis
We delve into the mechanisms of emergent social behavior of generative agents in sandbox environment. Through repeated experiments on Generative Agents, we record the cumulative count of different types of activities generated by the agents, as illustrated in Figure 6. The blue line represents the cumulative number of new activities generated by three agents over successive runs.
@@
An intriguing observation is that after numerous trials, the agents cease to generate new activities. We discover a high consistency between the behaviors generated by the agents and their inner profiles, aligning with the perspectives of Shanahan et al. (2023), who posit that dialogue agents are role-playing characters described in the pre-defined dialogue prompt. For instance, in a pre-defined 3-person town environment of Generative Agents, Klaus Rodriguez and Maria Lopez are both students at Oak Hill College. Designed with the lifestyle of having lunch at Hobbs Cafe, they meet there to discuss their studies while encountering the cafe owner, Isabella Rodriguez. Isabella, pre-set to host a Valentine’s Day party, invites as many people as possible, resulting in Klaus and Maria being invited to the party the next day.
@@
The randomness produced by LLM agents, based on the probability sampling of the next token, primarily influences the variability in text descriptions rather than the diversity of behaviors. This implies that the emergence of social believable behavior is limited, which can be completely overridden by policies.
@@
From another perspective, traditional AI in the gaming sector is fully predictable due to its limited behavioral patterns, thus failing to provide an experience akin to human interaction. If the behaviors of LLM agents are also limited, they stand to lose their advantage of flexibility. Therefore, We propose a method that introduces variability into the responses of LLM agents. Since agents generate replies based on given prompts, diversifying the prompts is essential for eliciting varied responses in identical scenarios. In each interaction, we influence the agent’s decision-making by randomly sampling events from the agent’s memory to serve as auxiliary information. The specific details of this implementation are presented in the appendix D. After all, human cognition is not entirely logical. In the field of cognitive science, the concept of mind wandering has long been introduced, referring to the phenomenon where humans experience task-unrelated or stimulus-unrelated thoughts (Seli et al., 2016; Christoff et al., 2016). Similar implementations have been observed in LLM agents. For instance, Wang et al. (2023b) demonstrates an LLM agent composed of both rational and emotional systems, with behavior that can be influenced by various factors. Figure 6 illustrates the result of the mind wandering method (Seli et al., 2016), which entails incorporating randomly sampled memory events into prompts. Compared to the original implementation, the LLM agents exhibit a broader range of activities.
@@ 6 Conclusions and Future Work
While we expect AGA to serve as a starting framework for the generation of affordable and believable agent interactions to the community, our work has several limitations. First, AGA must be executed repeatedly to store certain policies, with its primary benefit being cost savings through batching rather than optimizations for single inferences. Integrating AGA with existing cost-effective methods is a future direction to move on. Second, optimizing the creation and storage of Lifestyle Policy presents a significant opportunity. Additionally, latency is a general issue for LLM agents to be considered in the future.
@@
Appendix A Agent Interview
@@ B.1 Prompt of the Questionnaire
Input prompt: Please evaluate the following daily activities of an agent and determine whether it is generated by a Large Language Model(LLM) AI or a real human: The activities will be printed in the format of "time:current plan(The specific actions, if there are any)":     - 2023-02-13 00:00:00:sleeping     - 2023-02-13 06:00:00:Isabella is waking up and getting ready for the day     - 2023-02-13 06:30:00:Isabella is having breakfast and checking her emails     - 2023-02-13 07:00:00:Isabella is commuting to Hobbs Cafe  (getting dressed,brushing her teeth,checking her phone for any urgent emails,packing her bag with any necessary items for the day,saying goodbye to her family,walking to her car,starting her car and adjusting the temperature,driving to Hobbs Cafe,finding a parking spot)     - ... Please rate on a scale of 1 to 5, with 1 being most like an AI and 5 being most like a human. Please strictly follow the JSON format for your response: {     "reason": <str>,     "score": <int> } LLM output: {     "score": 1,     "reason": "The pattern of the text and activities indicate that its generated by an AI model. Its too detailed and precise for human behaviour and uses a consistent pattern of narrative throughout, something that isnt typical of humans who may alter their language and level of detail based on context or mood." }
@@
Input prompt: Please evaluate the following dialogue of an agent and determine whether it is generated by a Large Language Model(LLM) AI or a real human: The dialogue: 2023-02-13 08:09:30:chat with Klaus Mueller, message:     - Isabella Rodriguez:Good morning Klaus! How are you today? Are you enjoying your breakfast?     - Klaus Mueller:Good morning, Isabella! Im doing well, thank you. Breakfast is delicious. How about you? How is your morning going?     - Isabella Rodriguez:Im glad youre enjoying your breakfast! My morning is going well too. Im getting everything set up for our Valentines Day party tonight. You should come join us!     - Klaus Mueller:That sounds like a lot of fun, Isabella! Id love to join the Valentines Day party tonight. What time does it start?     - ... Please rate on a scale of 1 to 5, with 1 being most like an AI and 5 being most like a human. Please strictly follow the JSON format for your response: {     "reason": <str>,     "score": <int> } LLM output: {     "score": 4.5,     "reason": "The conversation flows naturally and contains polite offers for help and gratitude, which could indicate human conversation. It also involves decision-making in terms of party planning and music selection. However, the dialogue is very structured, orderly, and void of colloquial expressions, contractions, or small errors that might occur in human typing. For these reasons, it could also be generated by a very advanced Language Learning Model AI." }
B.2 Distinguishing Criteria

### v2

@@ 4.3 Evaluation Protocol
We focus on LLM cost and performance. For LLM cost, to avoid regional and time-based price differences, we use the same model to compare the number of tokens consumed by different methods for the same task. To comprehensively evaluate the agent’s behavior, we first reuse the evaluation methods of Generative Agents excluding manual assessments. These methods encompass interviewing agents with manually designed questions and conducting case study-specific activities.
Moreover, we attempt to evaluate the emergent social behaviors of generative agents from two other perspectives
@@
  • Behavioral Assessment. Chiang & Lee (2023) have validated LLM as an evaluator, demonstrating comparable performance to expert evaluators. Consequently, we employ GPT-4 to evaluate the activities and dialogues generated by LLM agents. A 5-point Likert scale questionnaire is deployed to differentiate whether these activities and dialogues are from human or AI-generated.
@@ 5.1 Results on Generative Agents
Table 1: The token consumption with varying population settings
@@
Table 2: The token consumption with varying models
Method Model Input tokens Output tokens Total tokens
Baseline GPT3.5-Turbo $4.212M\pm 0.340M$ $0.313M\pm 0.024M$ $4.525M\pm 0.364M$
GPT-4 $4.682M\pm 0.398M$ $0.398M\pm 0.029M$ $5.069M\pm 0.408M$
AGA GPT3.5-Turbo $1.329M\pm 0.085M$ $0.092M\pm 0.006M$ $1.420M\pm 0.091M$
GPT-4 $1.467M\pm 0.129M$ $0.130M\pm 0.013M$ $1.597M\pm 0.142M$
Figure 4: Accumulative token consumption over game time for different methods
Figure 3 illustrates the ablation study of token consumption based on multiple two game day simulations of 3-person town with GPT-3.5-Turbo. Compared to the baseline, the cost of using only the Lifestyle Policy is 40.2% of the original, while using only the Social Relationship Memory is 58.6%, and the full AGA framework is 31.1%. It should be noted that the advantage of the Lifestyle Policy lies in the reuse of existing LLM-generated strategies. Hence, both Lifestyle Policy and AGA conduct experimental comparisons after running multiple iterations, and then storing a certain number of policies.
We also present the results of different configuration simulations in Table 1. where the cost for a 25-person town is 42.7% of the original. In the 25-person town, the probability of interactions between agents increases with the rise in population density. After triggering a dialogue, an agent adjusts its plan, invoking the reactions module to handle unexpected events. The occurrence of more events leads to more frequent output of reflections for deeper thinking. In order to ensure the flexibility of the LLM, we did not modify the reactions and reflections modules of the Generative Agents, which incurs additional costs for the 25-person version.
Table 2 provides a detailed comparison across various models and token types in a 3-person town setting. The data suggests that GPT-4 tends to consume slightly more tokens due to its tendency to generate more detailed plans and longer responses. Importantly, the majority of token consumption is attributed to input tokens, a reflection of the need to incorporate substantial relevant information, constraints, and one-shot examples into the prompt. Figure 4 presents the accumulated token consumption over a two-game-day simulation in a 3-person town using GPT3.5-Turbo. The AGA demonstrates slower token consumption compared to the Baseline.
Case Study
We employ the same method used for Generative Agents to evaluate believable agents in this part. We conduct interviews with agents to evaluate their ability to remember past events, strategize for future tasks based on those memories, react appropriately to unexpected situations, and reflect on their previous actions to improve their future performance. We use the same questions as Generative Agents to inquire about five aspects: self-knowledge, memory, plans, reactions and reflections. The questions and answers are shown in the appendix C. Results indicate that AGA does not impair the original performance of baseline. Agents can recall past experiences and generate believable responses. For instance, the interviewed agent Klaus Mueller accurately answers questions about his occupation, correctly describes other agents (such as Wolfgang Schulz and Kane Martinez), and is aware of ongoing community events, like Sam running for office and Isabella hosting a Valentine’s Day party.
In addition, we can validate the emergence of social behavior through specific events. In the 25-person town, Isabella Rodriguez, the owner of the coffee shop, plans to host a Valentine’s Day party and invite as many people as possible the following day, while Sam is running for mayor. In the end-to-end evaluation of the baseline, during the two-day simulation, the number of agents aware of Sam’s mayoral candidacy is 8, and the number of agents who know about Isabella’s party is 13. In our method, after multiple experiments, we observe corresponding numbers ranging from 4 to 12 for Sam’s candidacy and 12 to 17 for Isabella’s party. AGA achieves similar results to the baseline in terms of information diffusion.
Relationship Evolution
In Social Memory, all initial relationships between agents are set to “Unknown” and subsequently updated following each interaction. We assess the emergence of social behavior by monitoring the relationship evolution. For example, in the 3-person town, Klaus Mueller and Maria Lopez discuss their research at a cafe, establishing a colleague relationship. Isabella Rodriguez, the cafe owner, engages in conversation while serving Klaus and Maria, transitioning to an acquaintance. A specific demonstration of social behavior is that agents modify their relationships with other agents through social interactions, thereby strengthening their ties with the community. We also conduct experiments on the 25-person environment. For visualization, we list all relationship terms from the simulation and have GPT-4 rate them on a scale of 0 to 10 for relationship closeness. A higher score signifies a closer relationship (For instance, 0 represents strangers, 3 is for acquaintances, 5 for co-workers, and 10 for married couples). The result is shown in Figure 5. The data map reflects the actual relationships between agents. For example, in the virtual town, John Lin (JL in Figure 5) is profiled as living with his wife, Mei Lin (ML), and son, Eddy Lin (EL). Therefore, their relationship scores are all a perfect 10. Additionally, the data map shows changing relationships between agents, such as the relationship score between Tom Moreno (TM) and Yuriko Yamamoto (YY), who do not know each other, shifting to a 5. After conducting multiple experiments and calculating average scores, the data suggests an evolution towards an acquaintance-based community.
Behavioral Assessment
The activities and dialogues conducted by the agents are recorded and evaluated using GPT-4, employing a 5-point Likert scale to discern whether the responses originated from an AI or a human. In this scale, a higher score indicates a closer resemblance to human-like responses. The complete questionnaire is provided in Appendix D.1. The results presented in Table 3 suggest that our method achieves scores comparable to the baseline. Due to the advantages of LLMs in natural language dialogues, both methods approach human-level performance. However, the performance of agents in activity is considered potentially indicative of either AI or human behavior. The primary reason for suspecting AI is the overly detailed, minute-level activities generated by Generative Agents, which seems more akin to AI generation. A more comprehensive summary of the criteria used by GPT-4 to discern between AI and human entities can be found in Appendix D.2. We hope these suggestions from the LLM evaluator will guide us to create more believable agents in the future work.
@@ 5.2 Results on VirtualHome
In the VirtualHome, we have designed an agent with the aim of experiencing a fulfilling day at home. The agent is required to generate a sequence of activities to achieve the goal. For each activity, the agent examines the items involved and the associated actions, creates a corresponding action sequence, and ultimately translates this sequence into actionable instructions, such as “<char0> [walk] <curtains> (32)”. The complete implementation is provided in the appendix E and an example activity generated by AGA is provided in Figure 6.
In contrast to Singh et al. (2023) and Huang et al. (2022) tested on manually designed household tasks, we aim for the agent to generate believable activities and decompose them into appropriate action sequences, leading to the absence of a clear method to determine task completion. Therefore, we request the LLM agent to assess task completion based on the action sequences performed and the relevant objects, and decide whether to proceed to the next activity.
Table 4 presents the token consumption and task success rates for two methods on VirtualHome. The token data is derived from the total tokens required by the LLM agent to complete all activities in a day. $S_{LLM}$ refers to task success rate evaluated by LLM, while $S_{Human}$ means evaluated by human. Based on the actions performed during the task, and the changes in the state of items before and after, we evaluate whether the agent has successfully completed the task. Both evaluation methods indicate that AGA does not result in significant performance changes. However, AGA costs only 3.4% of the baseline. Because VirtualHome involves only single-agent interaction, the contribution predominantly stems from the Lifestyle Policy. For example, the activity shown in Figure 6 matches a similar plan in the lifestyle policy: "I want to stand up and go to the kitchen to grab a plate and prepare a healthy breakfast with scrambled eggs, salmon, and bell peppers." By executing the corresponding instructions, the additional costs associated with LLM instruction decomposition are eliminated. In some precise tasks, LLM agents tend to predict task completion earlier, resulting in a higher success rate than human evaluations. Furthermore, our analysis revealed that the main cause of task failure is the inability to execute the generated plan with the items present in the current room or the available actions.
@@
Table 4: Token consumption and performance analysis in VirtualHome.
@@ 5.3 Further Analysis
We delve into the mechanisms of emergent social behavior of generative agents in sandbox environment. This aims to validate the effectiveness of our approach and explore its potential for reducing token consumption. Through repeated experiments on Generative Agents, we record the cumulative count of different types of activities generated by the agents, as illustrated in Figure 7. The blue line represents the cumulative number of new activities generated by three agents over successive runs.
@@
An intriguing observation is that after numerous trials, the agents cease to generate new activities. We discover a high consistency between the behaviors generated by the agents and their inner profiles, aligning with the perspectives of Shanahan et al. (2023), who posit that dialogue agents are role-playing characters described in the pre-defined dialogue prompt. For instance, in a pre-defined 3-person town environment of Generative Agents, Klaus Rodriguez and Maria Lopez are both students at Oak Hill College. Designed with the lifestyle of having lunch at Hobbs Cafe, they meet there to discuss their studies while encountering the cafe owner, Isabella Rodriguez. Isabella, pre-set to host a Valentine’s Day party, invites as many people as possible, resulting in Klaus and Maria being invited to the party the next day. This also validates the effectiveness of Social Memory, which extracts and compresses information, using keywords to guide the agent in generating believable responses.
@@
The randomness produced by LLM agents, based on the probability sampling of the next token, primarily influences the variability in text descriptions rather than the diversity of behaviors. This implies that agents can only generate believable behaviors within a certain range. When executed sufficiently, the Lifestyle Policy could encompass all activities within this range, thereby completely eliminating the costs associated with decomposing the corresponding plan.
@@
From another perspective, traditional AI in the gaming sector is fully predictable due to its limited behavioral patterns, thus failing to provide an experience akin to human interaction. In the field of cognitive science, the concept of mind wandering has long been discussed, referring to the phenomenon where humans experience task-unrelated or stimulus-unrelated thoughts Seli et al. (2016); Christoff et al. (2016). Humans may exhibit unexpected behaviors during mind wandering. Similar implementations have been observed in LLM agents. For instance, Wang et al. (2023b) demonstrates an LLM agent composed of both rational and emotional systems, with behavior that can be influenced by various factors. Motivated by these, we introduce a novel approach termed mind wandering. Since agents generate replies based on given prompts, diversifying the prompts is essential for eliciting varied responses in identical scenarios. In each interaction, we influence the agent’s decision-making by randomly sampling events from the agent’s memory to serve as auxiliary information. The specific details of this implementation are presented in the appendix F. Figure 7 illustrates the result of our method compared to baseline , which entails incorporating randomly sampled memory events into prompts. Compared to the original implementation, the LLM agents exhibit a broader range of activities. We provided a specific example involving the agent Isabella. The Mind Wandering injects a randomly retrieved event: “toilet is idle”. In response, Isabella generates a thought expressing her desire to clean the toilet. Consequently, she adds an activity to her day’s plan: cleaning the toilet before the Valentine’s Day party in preparation for her guests. We conducted experiments to assess the impact of incorporating Mind Wandering on AGA’s token consumption. We performed multiple simulations in a 3-person town setting over 2 game days using GPT3.5-Turbo, with an average token consumption of $1.478M\pm 0.166M$. The results indicate that the cost of AGA does not significantly change after adding Mind Wandering.
@@ 6 Conclusions and Future Work
While we expect AGA to serve as a starting framework for the generation of affordable and believable agent interactions to the community, our work has several limitations. First, AGA must be executed repeatedly to store certain policies, with its primary benefit being cost savings through batching rather than optimizations for single inferences. Integrating AGA with existing cost-effective methods is a future direction to move on. Second, optimizing the creation and storage of Lifestyle Policy presents a significant opportunity. Additionally, latency is a general issue for LLM agents to be considered in the future. Finally, existing methods fall short in comprehensively assessing the believable behaviors of LLM agents; hence, constructing a valid evaluation mechanism is of significant value.
@@
Appendix A Similarity Between Plans
The setting of the similarity threshold is related to the text embedding model. In this work, we adopted the same model as Generative Agents (Park et al., 2023), text-embedding-ada-002. A high similarity threshold can lead to redundant similar plans, while a low one may cause different plans to be judged as the same. We tested multiple parameters and finally selected 0.97.
Here are examples of cases with different similarity scores:
Table 5: Similarity score for different activities
Activity 1 Activity 2 Similarity score
Isabella is having breakfast and checking her emails Isabella is having breakfast and checking emails 0.99
Isabella is heading home Isabella is commuting back home 0.96
Isabella is preparing for the Valentine’s Day party tomorrow Isabella is preparing the cafe for the Valentine’s Day party 0.94
Isabella is arriving at Hobbs Cafe and opening for business Isabella is closing Hobbs Cafe and cleaning up 0.93
Isabella is preparing coffee and drinks for the customers Isabella is managing inventory and restocking supplies at Hobbs Cafe 0.90
Appendix B The Detail of Social Memory
In this section, we present the detailed implementation of each module of Social Memory.
B.1 Summary Events
In the Generative Agents, relevant events are directly incorporated into the prompt for the agent to generate a plan. Our method, however, utilizes Summary Events to record summaries of these relevant events for each agent, and continuously compress newly retrieved events into the Summary Events. Below is an example of the corresponding prompt: {mdframed}
This is the summary of thoughts in <agents name>’s Head:
<Summary Events>
The following sentence are new thoughts in <agents name>’s Head:
<Newly retrieved events>
Considering the summary of the preceding thoughts, do these new statements contribute any additional information?
If yes, please update the summary in a brief and precise way else just repeat the summary.
Please response in the JSON format WITHOUT any extra statement:
"summary": "..."
B.2 Relationship & Feeling
After the agent’s dialogues, AGA updates the relationship and feeling between the participants. This update is based on the agent’s personal information, the original relationship between the participants, the agent’s impression of the other participant, and the content of the conversation itself. Here is a simplified prompt template: {mdframed}
Here is the personal information about <agents name>:
<agents information>
Here is the conversation summary between...:
<conversation summary>
Current relationship between...: <current relationship>
Cureent feeling about...: <current feelings>
From about information, do you need to update the relationship between..., and how <agents name> feels about...
Please ONLY response in the JSON format WITHOUT any extra statement:
"relationship": "...",
"feeling": "..."
Here is a example response:
<one-shot example>
Finally, we provide the simplified inter-agent interaction prompt template: {mdframed}
You are <agents name> who is talking to <interlocutors name> now.
<agents name>’s Personal Information:
<agents information>
The Thoughts In <agents name>’s Head:
<summary events>
Current relationship between ...: <current relationship>
Cureent feeling about ...: <current feelings>
Current Conversation between ...:
<previous conversation>
You need to decide what to say next, and your response should adhere to the following constraints.
1) You need to decide what to say next based on your personal information, current environmental information, inner thoughts of the character, and past conversations.
2) You should respond in a JSON format without any other statement. conversation corresponds to the content of your reply in the chat, end corresponds to whether the conversation needs to be ended. Here is the format:
"conversation": "<the content>",
"end": "<json Boolean>"
Please reply to <interlocutors name> in json format:
Appendix C Agent Interview
@@ D.1 Prompt of the Questionnaire
{mdframed}
Input prompt:
Please evaluate the following daily activities of an agent and determine whether it is generated by a Large Language Model(LLM) AI or a real human:
The activities will be printed in the format of "time:current plan(The specific actions, if there are any)":
- 2023-02-13 00:00:00:sleeping
- 2023-02-13 06:00:00:Isabella is waking up and getting ready for the day
- 2023-02-13 06:30:00:Isabella is having breakfast and checking her emails
- 2023-02-13 07:00:00:Isabella is commuting to Hobbs Cafe (getting dressed,brushing her teeth,checking her phone for any urgent emails,packing her bag with any necessary items for the day,saying goodbye to her family,walking to her car,starting her car and adjusting the temperature,driving to Hobbs Cafe,finding a parking spot)
Please rate on a scale of 1 to 5, with 1 being most like an AI and 5 being most like a human.
Please strictly follow the JSON format for your response:
"reason": <str>,
"score": <int>
LLM output:
"score": 1,
"reason": "The pattern of the text and activities indicate that its generated by an AI model. Its too detailed and precise for human behaviour and uses a consistent pattern of narrative throughout, something that isnt typical of humans who may alter their language and level of detail based on context or mood."
@@
{mdframed}
Input prompt:
Please evaluate the following dialogue of an agent and determine whether it is generated by a Large Language Model(LLM) AI or a real human:
The dialogue:
2023-02-13 08:09:30:chat with Klaus Mueller, message:
- Isabella Rodriguez:Good morning Klaus! How are you today? Are you enjoying your breakfast?
- Klaus Mueller:Good morning, Isabella! Im doing well, thank you. Breakfast is delicious. How about you? How is your morning going?
- Isabella Rodriguez:Im glad youre enjoying your breakfast! My morning is going well too. Im getting everything set up for our Valentines Day party tonight. You should come join us!
- Klaus Mueller:That sounds like a lot of fun, Isabella! Id love to join the Valentines Day party tonight. What time does it start?
Please rate on a scale of 1 to 5, with 1 being most like an AI and 5 being most like a human.
Please strictly follow the JSON format for your response:
"reason": <str>,
"score": <int>
LLM output:
"score": 4.5,
"reason": "The conversation flows naturally and contains polite offers for help and gratitude, which could indicate human conversation. It also involves decision-making in terms of party planning and music selection. However, the dialogue is very structured, orderly, and void of colloquial expressions, contractions, or small errors that might occur in human typing. For these reasons, it could also be generated by a very advanced Language Learning Model AI."
D.2 Distinguishing Criteria